# Simulation-Based Inference for Direction Reconstruction of Ultra-High-Energy Cosmic Rays with Radio Arrays

### v1

@@ II.3 Graph Neural Network Direction Estimator
Table 1: Prior distributions for the shower direction parameters used in the LtU-ILI training. Each Cartesian component of the unit direction vector is drawn independently from a uniform distribution on the interval [1,1][-1,1].
@@ II.4 Plane-Wave Front Prior as a Physics-Informed Constraint
To overcome limitations of purely data-driven neural networks, we incorporate a physics-informed prior based on the PWF approximation. This model analytically infers cosmic-ray arrival directions by fitting antenna timing patterns to a planar wavefront perpendicular to the shower propagation axis [26]. While neglecting curvature introduces systematic offsets—particularly for near-vertical tracks or impact points close to the array core—the model rigorously preserves causal ordering and global geometry. Our encoder follows the blueprint of Ref. [49], adapted to the present dataset and training strategy.
@@
$\mathbf{k}_{\text{final }}=\mathbf{k}_{\text{PWF }}+\alpha\,\Delta\mathbf{k},\quad\text{ with }\quad\left\|\mathbf{k}_{\text{final }}\right\|=1.$
@@ II.5 Simulation-based Inference within the LtU-ILI Framework
A training set of approximately 8,200 simulated air-shower events passing predefined quality criteria ( $\geq 5$ antennas triggered above a $60~\mu\mathrm{V}$ threshold) was used to train the neural posterior estimators. LtU-ILI manages model optimization by minimizing the negative log-posterior of the reconstructed cosmic-ray direction vectors, employing uniform priors to ensure unbiased generalization. Model training converges typically within 50 epochs on standard GPU hardware. Dataset splits (train/validation/calibration/test) and temperature-scaling calibration are described in Appendix A.
@@ III Results
Figure 4: Top: Joint posterior distribution for the mock GRANDproto300-like event shown in Fig. 3. Shaded regions represent the 68% and 95% credible intervals for zenith and azimuth, with the true arrival direction marked by a red cross. Corresponding one-dimensional marginal distributions for zenith (top axis) and azimuth (right axis) are also shown. Bottom: Percentile-percentile (P–P) plots for zenith (left) and azimuth (right), constructed from 1,560 test events. Empirical percentiles (blue) mostly follow the diagonal line and lie within the 95% bootstrap uncertainty (grey) band. A slight deviation above the diagonal at intermediate quantiles indicates mildly conservative posteriors.
@@
The network parameters are fitted exclusively on the training data; we tune learning-rate schedules and early-stopping thresholds on the remaining validation set and keep an untouched test set for the final metrics displayed in this section. Finally, we assessed the calibrated posterior using the coverage and sharpness diagnostics recommended by recent SBI benchmarks [50], ensuring that the reported credible regions are both compact and statistically well-calibrated.
Figure 5: TARP (Tests of Accuracy with Random Points) coverage plot for $1,560$ validation events (see Ref. [51]). The diagnostic was performed directly on the full direction vector $\mathbf{k}$, rather than separately on zenith and azimuth angles. The blue curve represents empirical expected coverage calculated from $2,048$ posterior samples per event, the shaded region indicates a bootstrap-derived $1\,\sigma$ uncertainty band, and the dashed diagonal marks perfect calibration. The observed near-diagonal alignment confirms well-calibrated posteriors, exhibiting mildly conservative uncertainty estimates at credibility levels above $40\%$.
@@
Figure 4 (bottom row) presents percentile–percentile (P–P) plots constructed from 2,048 independent validation events, comparing nominal posterior quantiles to their empirical coverage frequencies. The empirical curves closely follow the ideal one-to-one line, though a minor upward deviation at all quantiles suggests mildly conservative credible intervals, indicating slightly broader coverage than strictly required. This slight excess coverage is desirable in astrophysical analyses, where underestimating uncertainty carries a higher cost than modest conservatism [50].
To supplement the P-P coverage tests, we subjected our inferred posteriors to the Tests of Accuracy with Random Points (TARP) diagnostic introduced in Ref. [51]. TARP compares the nominal credibility $\alpha$ of many randomly-centered hyper-spherical regions to the empirical coverage probability estimated from posterior samples. Unlike traditional highest-posterior-density or simulation-based-calibration tests, TARP relies only on posterior samples and is both a necessary and a sufficient condition for calibration. In principle, the identity line is recovered if the approximate posterior equals the true one for every datum. Consequently, it can reveal subtle pathologies that other diagnostics may miss.
We ran the TARP coverage test in the LtU-ILI framework on the direction vectors $\mathbf{k}$. For each of the $1,560$ held-out events we drew 2,048 posterior samples from the flow ensemble. The expected-coverage curve hugs the one-to-one line across the full credibility range, drifting only slightly above it (mild conservatism) for credibility $\gtrsim 0.4$ (Fig. 5).
@@
Angular-resolution performance of the SBI pipeline
We measure directional precision with the Bayesian $68\%$ highest-posterior-density angle, $R_{68\%}$, defined as the radius of the smallest spherical cap that contains $68\%$ of the posterior mass for a single event. Because $R_{68\%}$ is rotation-invariant and tied directly to posterior mass, it avoids the coordinate artifacts that can distort marginal uncertainties in zenith and azimuth [52]. We compute $R_{68}$ for all 1,560 events in the hold-out validation split, representing each posterior with 2,048 Monte-Carlo samples; this depth keeps the Monte-Carlo noise in the quoted quantiles below $1\%$.
@@
We now examine whether the pipeline’s angular resolution varies systematically with event geometry. Figure 7 (left) reveals a clear zenith trend in the $68\%$ highest-posterior-density angle $R_{68\%}$. Resolution improves steadily from vertical showers to Earth-skimming events, with the tightest constraints at the most inclined geometries. The shaded band (median $\pm$ interquartile range) stays below $\approx 1^{\circ}$ across the full zenith span, showing that residual scatter arises mainly from measurement noise rather than zenith-dependent bias. We emphasize that we report median and interquartile range to capture the central tendency and spread without assuming Gaussian errors.
Impact of antenna multiplicity
@@
Our results suggest that adding more antennas will continue to narrow the posterior, but pushing the centroid error below the $\approx 0.3^{\circ}$ plateau would require sub-nanonsecond timing instrumentation. However, a complementary approach [49] would be to enrich our embedding network input by supplying per-station electic field amplitudes or energy fluence, which could allow the encoder to down-weight inconsistent arrival-amplitude combinations, thereby trimming the residual bias without requiring a hardware upgrade.
Figure 8: Timing-noise information limit. Median $68\%$ HPD angular radius $R_{68\%}$ is shown versus triggered-antenna multiplicity for $5\mathrm{~ns}$ jitter noise (solid blue, $16-84\%$ band) and no noise (orange markers, $16-84\%$ bars). Finer multiplicity bins than in Fig. 7 (right) clarify the high-multiplicity plateau that emerges once $\gtrsim$ 30 antennas are triggered. The residual panel ($R_{68\%}^{5\mathrm{~ns}}/R_{68\%}^{0\mathrm{~ns}}$ vs. multiplicity) shows that GPS jitter, not array geometry, sets the ultimate angular-resolution floor.
@@
Analytic direction reconstruction pipelines can reach very small point-estimate errors under certain conditions. In particular, for highly inclined showers ($\theta\gtrsim 60^{\circ}$) using timing and amplitudes, the Angular Distribution Function method [24] reports a mean direction error better than $4^{\prime}\left(\approx 0.07^{\circ}\right)$ on simulations. Furthermore, the PWF method [26], which uses timing only with a controlled timing model, provides closed-form error estimates and achieves $\approx 0.2^{\circ}$ mean errors at $10$ ns timing noise for very inclined events. By contrast, our results are timing-only and reported as calibrated Bayesian credible regions over $37^{\circ}-87^{\circ}$; the median $R_{68\%}$ is $\approx 0.38^{\circ}$ for $5$ ns clock jitter. These are not apples-to-apples conditions (inputs and metrics differ), but together they bracket the landscape: point estimator methods excel at sharp centroids, while our pipeline delivers calibrated, event-level uncertainty from sparse timing.
@@ IV Summary and Discussion
Our simulation-calibrated posteriors pass both PPP-P and TARP diagnostics, indicating that the quoted credible regions are not over-confident across the entire validation sample. Because calibration is performed with the same simulator that generated the training data, the pipeline implicitly marginalizes over all nuisance effects present in those simulations (e.g., geomagnetic variation, atmospheric refractivity, and radio-wave propagation uncertainties) without imposing parametric forms.
@@
For context, the analytic Angular Distribution Function method [8, 24] reports a sharper $\approx 0.1^{\circ}$ direction resolution under similar jitter noise. This is similar to the resolution found in the case of the plane wavefront reconstruction method discussed in Ref. [26]. However, it should be noted that their quoted uncertainties come from simulation ensembles rather than an event-level posterior. Our pipeline, therefore, complements these efforts by maintaining competitive resolution while adding verifiable, per-event credibility regions.
@@
Our encoder processes a four-dimensional $(x,y,z,t)$ input vector for each antenna, dynamically building a k-nearest-neighbour graph. This design remains independent of array geometry, allowing straightforward adaptation to IceCube-Gen2 [17], AugerPrime Radio [14], BEACON [15], or RNO-G [16] stations by retraining on the appropriate layouts. Additionally, since our residual-gated network combines an analytic plane-wave fit with data-driven corrections, other first-order wavefront approximations, such as spheroidal wavefronts or template libraries, could be integrated without substantial modifications.
@@ V Conclusions
We present the first end-to-end simulation-based inference pipeline for UHECR direction reconstruction using radio arrays that achieves a median $68\%$ containment of $0.38^{\circ}$ with correct coverage on realistic GRANDproto300 simulations.
@@
The training and inference code used in this work is archived on Zenodo (DOI: 10.5281/zenodo.16895985) and mirrored at GitHub. The training dataset used to produce the reported results consists of GRAND Collaboration proprietary ZHAireS simulations and cannot be redistributed. To facilitate reuse, the repository includes a lightweight synthetic/mock dataset and configuration files that allow users to run the full pipeline and explore the code. Trained model checkpoints derived from the proprietary data are not publicly released; requests for access to the underlying simulations should be directed to the corresponding author and are subject to GRAND Collaboration data policies.
@@ A.2 Graph Convolutional Network Architecture
We detail the embedding architecture in Table 2. Briefly, two graph convolutional layers (GCNConv) [53] transform the initial four-dimensional node features into a 64-dimensional representation, which is globally mean-pooled into a single event-level embedding. An MLP maps this embedding to a directional correction vector $\Delta\mathrm{k}_{\mathrm{GCN}}$. A parallel gating network takes as input the concatenation of $\Delta\mathbf{k}_{\mathrm{GCN}}$ and the analytic plane-wavefront (PWF) direction, outputting a scalar weight $\alpha$ that blends the corrections. The final corrected direction is normalised to unit length, ensuring physical validity. The full encoder and gating network comprise approximately $2.37\times 10^{4}$ trainable parameters.
@@ A.3 Training and Optimisation
which is a metric widely used in rotation-estimation tasks because it preserves Euclidean smoothness yet respects spherical geometry [54].
@@ Appendix B Masked autoregressive flow architecture and training
Masked Autoregressive Flows (MAFs) build a complex probability density by chaining invertible, autoregressive transformations whose triangular Jacobians allow closed-form log-determinants [55]. This property lets us compute the log-density of simulated parameters during training, while still operating in a likelihood-free setting for the data. MAFs therefore serve as efficient neural density estimators within our simulation-based Bayesian framework, combining expressive power with fast sampling at inference time [29].

### v2

@@ II.3 Graph Neural Network Direction Estimator
Table 1: Prior distributions for the shower direction parameters used in the LtU‑ILI training. Each Cartesian component of the unit direction vector is drawn independently from a uniform distribution on the interval [1,1][-1,1].
@@ II.4 Plane-Wave Front Prior as a Physics-Informed Constraint
To overcome limitations of purely data-driven neural networks, we incorporate a physics-informed prior based on the PWF approximation. This model analytically infers cosmic-ray arrival directions by fitting antenna timing patterns to a planar wavefront perpendicular to the shower propagation axis Ferrière et al. [2025]. While neglecting curvature introduces systematic offsets—particularly for near-vertical tracks or impact points close to the array core—the model rigorously preserves causal ordering and global geometry. Our encoder follows the blueprint of Ref. Ferrière and Benoit-Lévy [2025], adapted to the present dataset and training strategy.
@@
$\mathbf{k}_{\text{final }}=\mathbf{k}_{\text{PWF }}+\alpha\,\Delta\mathbf{k},\quad\text{ with }\quad\left\|\mathbf{k}_{\text{final }}\right\|=1.$ (1)
@@ II.5 Simulation-based Inference within the LtU-ILI Framework
A training set of approximately 8,200 simulated air-shower events passing predefined quality criteria ( $\geq 5$ antennas triggered above a $60~\mu\mathrm{V}$ threshold) was used to train the neural posterior estimators. LtU-ILI manages model optimization by minimizing the negative log-posterior of the reconstructed cosmic-ray direction vectors, employing uniform priors to ensure unbiased generalization. Model training converges typically within 50 epochs on standard GPU hardware. Dataset splits (train/validation/calibration/test) are described in Appendix A.
@@
Normalizing flows can be slightly miscalibrated even when accurate in location. We therefore apply a scalar temperature $\tau$ to the learned log-density and renormalize, $p_{\tau}(\bm{\theta}\mid\mathbf{x}_{\mathrm{obs}})\propto p(\bm{\theta}\mid\mathbf{x}_{\mathrm{obs}})^{1/\tau}$. We choose $\tau$ once on a held-out calibration split to match nominal and empirical credibility. Unless stated otherwise, all reported results use this calibrated posterior (details in App. B.2).
@@ III Results
Figure 4: Top: Joint posterior distribution for the mock GRANDproto300-like event shown in Fig. 3. Shaded regions represent the 68% and 95% credible intervals for zenith and azimuth, with the true arrival direction marked by a red cross. Corresponding one-dimensional marginal distributions for zenith (top axis) and azimuth (right axis) are also shown. Bottom: Percentile-percentile (P–P) plots for zenith (left) and azimuth (right), constructed from $2,048$ test events. Empirical percentiles (blue) lie mostly above the diagonal line and its 95% bootstrap uncertainty band, indicating mildly conservative posteriors.
@@
The network parameters are fitted exclusively on the training data; we tune learning-rate schedules and early-stopping thresholds on the remaining validation set and keep an untouched test set for the final metrics displayed in this section. Finally, we assessed the calibrated posterior using the coverage and sharpness diagnostics recommended by recent SBI benchmarks Hermans et al. [2022], ensuring that the reported credible regions are both compact and statistically well-calibrated.
Figure 5: TARP (Tests of Accuracy with Random Points) coverage plot for $1,560$ validation events (see Ref. Lemos et al. [2023]). The diagnostic was performed directly on the full direction vector $\mathbf{k}$, rather than separately on zenith and azimuth angles. The blue curve represents empirical expected coverage calculated from $2,048$ posterior samples per event, the shaded region indicates a bootstrap-derived $1\,\sigma$ uncertainty band, and the dashed diagonal marks perfect calibration. The observed near-diagonal alignment confirms well-calibrated posteriors, exhibiting mildly conservative uncertainty estimates at credibility levels above $40\%$.
@@
Figure 4 (bottom row) presents percentile–percentile (P–P) plots constructed from 2,048 independent validation events, comparing nominal posterior quantiles to their empirical coverage frequencies. The empirical curves closely follow the ideal one-to-one line, though a minor upward deviation at all quantiles suggests mildly conservative credible intervals, indicating slightly broader coverage than strictly required. This slight excess coverage is desirable in astrophysical analyses, where underestimating uncertainty carries a higher cost than modest conservatism Hermans et al. [2022].
To supplement the P-P coverage tests, we subjected our inferred posteriors to the Tests of Accuracy with Random Points (TARP) diagnostic introduced in Ref. Lemos et al. [2023]. TARP compares the nominal credibility $\alpha$ of many randomly-centered hyper-spherical regions to the empirical coverage probability estimated from posterior samples. Unlike traditional highest-posterior-density or simulation-based-calibration tests, TARP relies only on posterior samples and is both a necessary and a sufficient condition for calibration. In principle, the identity line is recovered if the approximate posterior equals the true one for every datum. Consequently, it can reveal subtle pathologies that other diagnostics may miss.
We ran the TARP coverage test in the LtU-ILI framework on the direction vectors $\mathbf{k}$. For each of the $1,560$ held-out events we drew 2,048 posterior samples from the flow ensemble. The expected-coverage curve hugs the one-to-one line across the full credibility range, drifting only slightly above it (mildly conservative) for credibility $\gtrsim 0.4$ (Fig. 5). In particular, the 2,048 posterior samples were drawn independently from the trained MAF posterior for a fixed mock event $\mathbf{x}_{\mathrm{obs}}$ (we sample $z\sim\mathcal{N}(0,I)$ and map $z$ to $\bm{\theta}=f^{-1}(z;\,\text{embedding}(\mathbf{x}_{\mathrm{obs}}))$) with no re-simulation of timing jitter at inference (see Ref. Papamakarios et al. [2018] for MAF sampling details). Our MAF flow model is described in App. B.
@@
Angular‐resolution performance of the SBI pipeline
We measure directional precision with the Bayesian $68\%$ highest-posterior-density angle, $R_{68\%}$, defined as the radius of the smallest spherical cap that contains $68\%$ of the posterior mass for a single event. Because $R_{68\%}$ is rotation-invariant and tied directly to posterior mass, it avoids the coordinate artifacts that can distort marginal uncertainties in zenith and azimuth Bonifazi [2009]. We compute $R_{68}$ for all 1,560 events in the hold-out validation split, representing each posterior with 2,048 Monte-Carlo samples; this depth keeps the Monte-Carlo noise in the quoted quantiles below $1\%$.
@@
We now examine whether the pipeline’s angular resolution varies systematically with event geometry. Figure 7 (left) reveals a clear zenith trend in the $68\%$ highest-posterior-density angle $R_{68\%}$. Resolution improves steadily from vertical showers to near horizontal events, with the tightest constraints at the most inclined geometries. The shaded band (median $\pm$ interquartile range) stays below $\approx 1^{\circ}$ across the full zenith span, showing that residual scatter arises mainly from measurement noise rather than zenith-dependent bias. We emphasize that we report median and interquartile range to capture the central tendency and spread without assuming Gaussian errors.
Impact of number of triggered antennas
@@
Our results suggest that adding more antennas will continue to narrow the posterior, but pushing the centroid error below the $\approx 0.3^{\circ}$ plateau would require sub-nanonsecond timing instrumentation. However, a complementary approach Ferrière and Benoit-Lévy [2025] would be to enrich our embedding network input by supplying per-station electic field amplitudes or energy fluence, which could allow the encoder to down-weight inconsistent arrival-amplitude combinations, thereby trimming the residual bias without requiring a hardware upgrade.
Figure 8: Timing-noise information limit. Median $68\%$ highest-posterior-density (HPD) angular radius $R_{68\%}$ is shown versus triggered-antenna multiplicity for $5\mathrm{~ns}$ jitter noise (solid blue, $16-84\%$ band) and no noise (orange markers, $16-84\%$ bars). Finer multiplicity bins than in Fig. 7 (right) clarify the high-multiplicity plateau that emerges once $\gtrsim$ 30 antennas are triggered. The residual panel ($R_{68\%}^{5\mathrm{~ns}}/R_{68\%}^{0\mathrm{~ns}}$ vs. multiplicity) shows that GPS jitter, not array geometry, sets the ultimate angular-resolution floor.
@@
Figure 9: Comparison of the point-estimate angular resolution of the PWF-prior fed into the network (outline histogram) and our trained SBI model (filled histogram). The angular error is computed as $\psi=\arccos(\hat{\bm{k}}\cdot\bm{k}_{\rm true})$, where $\hat{\bm{k}}$ and $\bm{k}_{\rm true}$ are the reconstructed and true direction vectors, respectively. The vertical dashed lines show the mean ($\mu_{\rm PWF}=0.40^{\circ}$, $\mu_{\rm SBI}=0.43^{\circ}$) for both methods evaluated on the same held-out test events.
Analytic direction reconstruction pipelines can reach very small point-estimate errors under certain conditions. In particular, for highly inclined showers ($\theta\gtrsim 60^{\circ}$) using timing and amplitudes, the Angular Distribution Function method Guelfand et al. [2025] reports a mean direction error better than $4^{\prime}\left(\approx 0.07^{\circ}\right)$ on mock observations. Similarly, the PWF method introduced in Ref. Ferrière et al. [2025], which uses trigger times and antenna positions, reports a mean error of approximately $0.2^{\circ}$ for a GPS jitter of $10$ ns. These quoted resolutions were obtained on simulation datasets that differ from the one used here (so absolute values are not directly comparable).
We observe that we do not aim to outperform PWF in point-estimate resolution in this work; our focus is calibrated uncertainty. Nevertheless, it is instructive to compare the direction-reconstruction resolution of the PWF Ferrière et al. [2025] and our amortized SBI model on equal footing. For this, we evaluate both models on the same events from the held-out test split. For the PWF, we use the standalone public package (PWF_reconstruction  [^2] https://github.com/arsenefer/PWF_reconstruction).
Figure 9 shows histograms of the point estimates obtained with both methods. As can be seen, our amortized SBI model has learned to defer to the PWF-prior for virtually all events. However, our SBI model does learn the event-by-event uncertainty around the PWF prior. We expand on this in the Appendix, where we assess the direction uncertainty calibration by comparing empirical coverage to target credibility.
@@ IV Summary and Discussion
Our simulation-calibrated posteriors pass both PPP-P and TARP diagnostics, indicating that the quoted credible regions are not over-confident across the entire validation sample. Because calibration is performed with the same simulator that generated the training data, the pipeline implicitly marginalizes over all nuisance effects present in those simulations (e.g., geomagnetic variation, atmospheric refractivity) without imposing parametric forms.
@@
For context, the analytic Angular Distribution Function method Decoene et al. [2021], Guelfand et al. [2025] reports a sharper $\approx 0.1^{\circ}$ direction resolution under similar jitter noise. This is similar to the resolution found in the case of the plane wavefront reconstruction method discussed in Ref. Ferrière et al. [2025]. However, it should be noted that Ref. Ferrière et al. [2025] provides event-dependent analytic Gaussian uncertainty ellipses and validates them with simulation ensembles; it does not construct an event-level Bayesian posterior. Our pipeline complements this by delivering calibrated Bayesian posteriors with per-event highest-posterior-density credibility regions.
@@
Our encoder processes a four-dimensional $(x,y,z,t)$ input vector for each antenna, dynamically building a k-nearest-neighbour graph. Though our architecture does not hard-code a specific antenna array layout, adapting our model to other layouts such as AugerPrime Radio Aab and others [2016], or BEACON Southall and others [2023], could be achieved by retraining on the target layouts. Our current simulations include contained events only; performance on uncontained cores or markedly different geometries must be established by retraining on representative simulations and revalidating the posterior calibration. Additionally, since our residual-gated network combines an analytic plane-wave fit with data-driven corrections, other first-order wavefront approximations (e.g., spheroidal or hyperbolic wavefronts, or template libraries) could be integrated without substantial modifications.
@@ V Conclusions
We present the first end-to-end simulation-based inference pipeline for UHECR direction reconstruction using radio arrays that achieves a median $68\%$ containment of $0.38^{\circ}$ with correct coverage on realistic GRANDproto300 simulations. While our SBI model can represent non-Gaussian posteriors, the main advantage demonstrated here is not exotic posterior shapes, but rather validated calibration of credible regions under the full simulation procedure used to generate the training data.
@@
The training and inference code used in this work is archived on Zenodo Macias [2025] (DOI: 10.5281/zenodo.16895985) and mirrored at GitHub. The training dataset used to produce the reported results consists of GRAND Collaboration proprietary ZHAireS simulations and cannot be redistributed. To facilitate reuse, the repository includes a lightweight synthetic/mock dataset and configuration files that allow users to run the full pipeline and explore the code. Trained model checkpoints derived from the proprietary data are not publicly released; requests for access to the underlying simulations should be directed to the corresponding author and are subject to GRAND Collaboration data policies.
@@ A.2 Graph Convolutional Network Architecture
We detail the embedding architecture in Table 2. Briefly, two graph convolutional layers (GCNConv) Kipf and Welling [2017] transform the initial four-dimensional node features into a 64-dimensional representation, which is globally mean-pooled into a single event-level embedding. An MLP maps this embedding to a directional correction vector $\Delta\mathrm{k}_{\mathrm{GCN}}$. A parallel gating network takes as input the concatenation of $\Delta\mathbf{k}_{\mathrm{GCN}}$ and the analytic plane-wavefront (PWF) direction, outputting a scalar weight $\alpha$ that blends the corrections. The final corrected direction is normalised to unit length, ensuring physical validity. The full encoder and gating network comprise approximately $2.37\times 10^{4}$ trainable parameters.
@@ A.3 Training and Optimisation
which is a metric widely used in rotation-estimation tasks because it preserves Euclidean smoothness yet respects spherical geometry Geist et al. [2024].
@@ Appendix B Masked autoregressive flow architecture and training
Masked Autoregressive Flows (MAFs) build a complex probability density by chaining invertible, autoregressive transformations whose triangular Jacobians allow closed-form log-determinants Papamakarios et al. [2018]. This property lets us compute the log-density of simulated parameters during training, while still operating in a likelihood-free setting for the data. MAFs therefore serve as efficient neural density estimators within our simulation-based Bayesian framework, combining expressive power with fast sampling at inference time Ho and others [2024].
@@
For a given event $x$, we compute its 64-dimensional embedding $h(x)$ with the frozen encoder (the encoder blends the plane-wavefront seed $\mathbf{k}_{\mathrm{PWF}}$ with the learned residual $\Delta\mathbf{k}$ via a scalar gate $\alpha$ ). To draw one posterior sample, we draw $z\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}_{3}\right)$ and apply the inverse autoregressive transformation conditioned on $h(x)$ to obtain $\mathbf{k}=f^{-1}(z\mid h(x))$. We then re-normalise $\mathbf{k}$ to unit length. Repeating this procedure yields i.i.d. samples from $p(\mathbf{k}\mid x)$, which we use to form highest-posterior-densities and to compute coverage.